### v2

# Can LLMs Hit Moving Targets? Tracking Evolving Signals in Corporate Disclosures

Abstract
Moving targets (Cohen & Nguyen, 2024) – managers’ strategic shifting of key performance metrics when the original targets become difficult to achieve – have been shown to predict subsequent stock underperformance. However, our work reveals that the method employed in that study exhibits two key limitations that hinder the accuracy – noise in the extracted targets and loss of contextual information – both of which stem primarily from the use of a named entity recognition (NER). To address these two limitations, we propose an LLM-based target extraction method with a newly defined metric that better captures semantic context. This approach preserves semantic context beyond simple entity recognition and yields consistently higher predictive power than the original approach. Overall, our approach enhances the granularity and accuracy of financial text-based performance prediction.
Machine Learning, ICML
1 Introduction
Recent advances in Large Language Models (LLMs) have spurred efforts to leverage these tools for uncovering economically meaningful signals from unstructured financial texts such as earnings calls and corporate disclosures. In line with prior research, our work applies LLMs to enhance the target and metric extraction methodology proposed in Cohen & Nguyen (2024). Their study documents that managers strategically shift emphasis to new performance metrics when existing ones deteriorate – a “moving targets” behavior that predicts subsequent stock price declines.
Unlike the original paper, which relied on a Named Entity Recognition (NER)-based approach to extract and quantify moving targets, this study introduces an LLM-enhanced methodology. We argue that leveraging the contextual understanding of LLMs enables more interpretable and accurate measurement of strategic communication patterns in financial disclosures.
Through the rest of the paper, we introduce and validate our methodology and its effectiveness. Our contributions can be summarized as follow:
  • In Sec. 2.1, we identify two key limitations of the original NER-based methodology – noise in the extracted targets and loss of contextual information – that hinder accurate measurement of moving targets.
  • In Sec. 2.2, we propose LLM-driven solution to overcome these limitations: an LLM-based target extraction method with a newly defined metric based on semantic-similarity.
  • Our experiment in Sec. 3 demonstrates that these LLM-based enhancements deliver more precise, interpretable, and predictive measures of firms’ moving target behavior.
Figure 1: An illustration of the metric computation for LLM-based extraction leveraging semantic similarity. First, we extract targets from the firm’s earnings call transcripts at year-quarter $t$ and at t4t-4 using a LLM. Second, all extracted targets are encoded as vector embeddings using text encoder $f_{T}(\cdot)$. Third, for each target at t4t-4, we compute pairwise cosine similarities with targets at $t$, apply max-pooling to identify the most similar match, and perform thresholding $g(\cdot)$ to correct for paraphrases and remain robust to wording or formatting variations. Finally, we average the retained similarities to obtain the firm $c$’s Moving Targets score at time $t$, denoted $MT_{c,t}$.
2 Methodology
In this section, we present LLM-driven methods that enhance the measurement of firms’ moving targets from earnings call transcripts by overcoming key limitations of the original NER-based approach. We first motivate our method by pinpointing two potential sources of inaccuracy—noise in extracted targets and loss of contextual information. We then introduce an LLM-based target extraction method that computes a moving targets score leveraging semantic-similarity. The overall filtering and metric-computation processes are illustrated in Fig. 1.
2.1 Motivation
Prior work (Cohen & Nguyen, 2024) has relied heavily on NER to extract performance targets from earnings call transcripts. However, we found that this approach suffers from limitations in both accuracy and explainability.
First, Fig. 3 shows that the NER-based method extracts generic words, indicating the potential problem of noise. Second, Fig. 4 shows that contextual qualifiers are dropped, highlighting the potential loss of contextual information.
Such noise and context loss may introduce distortions in measuring the strategic shifting of key performance metrics. As a result, NER-based extraction alone has limitations in precisely capturing firms’ strategic shifts in performance metrics. To address these issues, we propose an LLM-based target extraction and scoring method.
Notation. Let $c$ denote a firm and let $t$ denote a year–quarter index (e.g., 2025 Q3). The text of firm $c$’s earnings call transcript at year–quarter $t$ is denoted by $\mathbf{T}_{c,t}$. We define two finite sets of targets extracted from these transcripts as $\mathcal{T}_{c,t}=\{T^{1}_{c,t},\dots,T^{n}_{c,t}\}$ and $\mathcal{T}_{c,t-4}=\{T^{1}_{c,t-4},\dots,T^{m}_{c,t-4}\}$, where $T^{i}_{c,t}$ and $T^{j}_{c,t-4}$ denote the $i$-th and $j$-th extracted targets at times $t$ and t4t-4, respectively. For simplicity, we denote that $n=|\mathcal{T}_{c,t-4}|$ and $m=|\mathcal{T}_{c,t}|$ are the corresponding set sizes. We denote the LLM employed for post-filtering and extracting of targets as $h_{\text{LLM}}$.
2.2 LLM-Based Target Extraction
In this section, we describe our proposed LLM-based target extraction and formulation of the moving targets score that leverages semantic similarity between targets mentioned in a firm’s earnings call transcripts. Fig. 1 provides an overview of the full computation pipeline. In the remainder of this subsection we formalize the notation and describe each component to compute $MT_{c,t}$, which represents the moving targets score of firm $c$ at year-quarter $t$.
Target Extraction. For each firm $c$ and year–quarter $t$, we process the entire transcript $\mathbf{T}_{c,t}$ in a single LLM call, with a prompt $P_{\text{ext}}$ that instruct the model $h_{\text{LLM}}$ to separately extract all performance targets from the management presentation and from the analyst Q&A sections. This procedure $h_{\text{LLM}}(T^{i}_{c,t}|P_{\text{ext}})$ also yields a finite set of targets $\mathcal{T}_{c,t}$ for firm $c$ in year-quarter $t$. This LLM-based extraction replaces original NER-based extraction and is designed to capture semantically rich targets that may be expressed through paraphrase or complex sentence structures.
Semantic Similarity. In the original formulation of the moving targets (Cohen & Nguyen, 2024), the degree of movement is quantified as
[…]
In particular, the numerator represents the number of missing targets, quantifying those previously emphasized performance metrics that no longer appear in the current transcript $\mathbf{T}_{c,t}$. Unlike the simple keyword matching used in NER, which allows straightforward counting, the LLM-based extraction often yields semantically equivalent targets expressed with different wording, making direct application of the original counting formula difficult.
To overcome this limitation, we introduce an semantic similarity-based metric. First, the extracted targets $T^{i}_{c,t}$ and $T^{j}_{c,t-4}$ is encoded into a dense vector embedding $\mathbf{e}(T^{i}_{c,t})=f_{T}(T^{i}_{c,t})$ using a off-the-shelve pre-trained text encoder $f_{T}(\cdot)$ that captures contextual meaning. We then compute all pairwise cosine similarities
$S_{ij}=\cos\bigl(\mathbf{e}(T^{i}_{c,t}),\;\mathbf{e}(T^{j}_{c,t-4})\bigr)$ (1)
between current and previous targets. For each previous target $T^{j}_{c,t-4}$, we apply max-pooling
$\tilde{S}_{j}=\max_{i}S_{ij}$
to identify its best semantic match in the current quarter.
Pooling. Suppose there exist two targets that convey the same meaning but differ slightly in wording across the calls at $t$ and t4t\!-\!4. If we rely on max-pooling alone, such pairs may receive lower similarity scores, causing them to be penalized as if they were different targets. To address this issue, we apply a thresholding function that treats semantically equivalent targets as fully retained:
$g(\tilde{S}_{j})=\begin{cases}1,&\tilde{S}_{j}\geq\tau,\\ \tilde{S}_{j},&\tilde{S}_{j}<\tau,\end{cases}$ (2)
with cutoff $\tau$ so that semantically equivalent targets are counted as fully retained, thereby correcting for paraphrases and making the metric robust to minor wording or formatting changes.
Scoring. Finally, we define the Moving Targets score as
$MT_{c,t}=\frac{1}{n}\sum_{j=1}^{b}g(\tilde{S}_{j}).$
This semantic similarity–based metric extends the original discrete measure by preserving semantic equivalence even when targets are phrased differently. This metric quantifies the extent to which performance targets previously emphasized by a firm are dropped or semantically altered after four quarters, similar with Cohen & Nguyen (2024).
3 Experiments
In this section, we empirically evaluate the effectiveness of our proposed LLM-based method. We begin by describing the dataset and summary statistics, then outline the implementation details and validation experiments. Finally, we present results that assess whether the LLM-based method reduces noise, preserves contextual information, and improves the predictive accuracy over NER-based method.
3.1 Data and Summary Statistics
We conduct our analysis using earnings call transcripts as the primary source of firm-level disclosures, following Cohen & Nguyen (2024). While we also experimented with 10-K and 10-Q MD&A sections, we found that earnings calls provided the most informative setting for detecting shifting in targets. Our study focuses on a restricted sample of S&P 100 firms due to computational cost. The sample covers between January 2010 and December 2024, producing 5,615 firm-quarter observations across 64 quarters.
On average, the LLM-based method extracted 150.21 targets per call, with 77.95 originating from the presentation and 76.89 from the Q&A. Of the 5,176 observations for which a moving target score could be computed, the mean score was 0.8713, with a standard deviation of 0.0848. On the other hand, the NER-based method extracted 40.36 targets per call, with 31.77 originating from the presentation and 9.61 from the Q&A. The mean score was 0.7147, with a standard deviation of 0.1656.
3.2 Implementation Details
To extract performance targets, we employ Google’s Gemini-2.5-Pro model and design prompts that separately request targets from the management presentation and the analyst Q&A portions of each transcript, as shown in Tab. 5. This design enables the model $h_{\text{LLM}}$ to identify semantically rich and context-aware targets that may be expressed through paraphrase, in contrast to the keyword-oriented NER extraction. Additionally, we employ OpenAI’s text-embedding-3-large model as an off-the-shelf pre-trained text encoder $f_{T}(\cdot)$ in Eq. 1. We set $\tau=0.65$ for the threshold in Eq. 2.
For validation, we adopt the empirical design of Cohen & Nguyen (2024). We construct equal-weighted quintile portfolios sorted on the moving targets score in a calendar-time framework, and we estimate Fama–MacBeth cross-sectional regressions with standard controls. This design allows us to test whether the LLM-based scores retain predictive power for subsequent stock returns, thereby providing a direct comparison to the original methodology.
Figure 2: Comparison of excess returns, Fama–French 3-factor alphas, and 5-factor alphas under the NER-based vs. LLM-based scoring approaches. Estimates are based on a calendar-time portfolio strategy that goes long in firms in the lowest quintile (Q1) of the Moving Targets distribution and short in those in the highest quintile (Q5). This design captures the return spread between firms with relatively fewer abandoned targets and those with more pronounced target shifting. The LLM-based approach consistently delivers stronger predictive power than the NER-based one.
3.3 Results
The LLM-based method extracts targets with reduced noise. Recall that our analysis shows the NER-based method frequently extracts generic fragments that lack clear economic meaning, as presented in Tab. 1. In contrast, Tab. 2 show that the LLM-based method identifies substantively meaningful items such as market share, revenue growth, cash flow, and dividends. This indicates that LLM-based extraction substantially mitigates the noise problem and more accurately captures firms’ moving target metrics.
Targets extracted by the LLM-based method retain richer contextual information. Another limitation of NER-based extraction is the systematic loss of contextual qualifiers. For example, as shown in Fig. 4, NER extracts only broad terms like revenue, dropping distinctions across segments such as Data Center revenue or Compute revenue. In contrast, Fig. 5 illustrates that the LLM-based method preserves these qualifiers, producing more granular targets such as Quarterly Data Center revenue growth or Quarterly Compute revenue growth. This richer contextual information enhances explainability and allows a more precise mapping between managerial communication and firm fundamentals.
The LLM-based method enhances predictive accuracy for subsequent stock performance declines. We next evaluate the economic value of the extracted metrics using portfolio sorts and Fama–MacBeth regressions. Tab. 3 shows that quintile portfolios formed on the LLM-based Moving Targets score exhibit significantly negative Q5–Q1 spreads, with alphas remaining robust after controlling for standard risk factors. By contrast, the NER-based scores yield weaker and mostly insignificant spreads. Consistent evidence from Fig. 2 and Tab. 4 indicates that the LLM-based score loads with stronger negative coefficients in cross-sectional regressions of returns, offering predictive power for subsequent stock underperformance. Together, these results confirm that LLM-based extraction not only reduces noise and preserves context, but also delivers a more economically meaningful predictor of firms’ future stock performance.
4 Conclusion & Future Work
This paper proposed an LLM-based method for measuring firms’ moving targets from earnings call transcripts. By reducing noise, preserving contextual detail, and improving predictive accuracy, our approach extends the original NER-based method and delivers more economically meaningful estimates of firms’ target shifting.
Future work may further enhance this method by analyzing sector-specific variations in moving targets. In particular, industry- and sector-level classification, combined with contextual prompt engineering, could yield even more precise and explainable measures of managerial emphasis across heterogeneous business environments.

### v3

@@
 Abstract
-Moving targets (Cohen & Nguyen, 2024) – managers’ strategic shifting of key performance metrics when the original targets become difficult to achieve – have been shown to predict subsequent stock underperformance.
+Moving targets (cohen2024moving) – managers’ strategic shifting of key performance metrics when the original targets become difficult to achieve – have been shown to predict subsequent stock underperformance.
 However, our work reveals that the method employed in that study exhibits two key limitations that hinder the accuracy – noise in the extracted targets and loss of contextual information – both of which stem primarily from the use of a named entity recognition (NER).
@@
 Recent advances in Large Language Models (LLMs) have spurred efforts to leverage these tools for uncovering economically meaningful signals from unstructured financial texts such as earnings calls and corporate disclosures.
-In line with prior research, our work applies LLMs to enhance the target and metric extraction methodology proposed in Cohen & Nguyen (2024).
+In line with prior research, our work applies LLMs to enhance the target and metric extraction methodology proposed in cohen2024moving.
 Their study documents that managers strategically shift emphasis to new performance metrics when existing ones deteriorate – a “moving targets” behavior that predicts subsequent stock price declines.
@@
 2.1 Motivation
-Prior work (Cohen & Nguyen, 2024) has relied heavily on NER to extract performance targets from earnings call transcripts.
+Prior work (cohen2024moving) has relied heavily on NER to extract performance targets from earnings call transcripts.
 However, we found that this approach suffers from limitations in both accuracy and explainability.
@@
 This LLM-based extraction replaces original NER-based extraction and is designed to capture semantically rich targets that may be expressed through paraphrase or complex sentence structures.
-Semantic Similarity. In the original formulation of the moving targets (Cohen & Nguyen, 2024), the degree of movement is quantified as
+Semantic Similarity. In the original formulation of the moving targets (cohen2024moving), the degree of movement is quantified as
 […]
@@
 Scoring. Finally, we define the Moving Targets score as
-$MT_{c,t}=\frac{1}{n}\sum_{j=1}^{b}g(\tilde{S}_{j}).$
+$MT_{c,t}=1-\frac{1}{n}\sum_{j=1}^{n}g(\tilde{S}_{j}).$
 This semantic similarity–based metric extends the original discrete measure by preserving semantic equivalence even when targets are phrased differently.
-This metric quantifies the extent to which performance targets previously emphasized by a firm are dropped or semantically altered after four quarters, similar with Cohen & Nguyen (2024).
+This metric quantifies the extent to which performance targets previously emphasized by a firm are dropped or semantically altered after four quarters, similar with cohen2024moving.
 3 Experiments
@@
 3.1 Data and Summary Statistics
-We conduct our analysis using earnings call transcripts as the primary source of firm-level disclosures, following Cohen & Nguyen (2024).
+We conduct our analysis using earnings call transcripts as the primary source of firm-level disclosures, following cohen2024moving.
 While we also experimented with 10-K and 10-Q MD&A sections, we found that earnings calls provided the most informative setting for detecting shifting in targets.
@@
 2.
-For validation, we adopt the empirical design of Cohen & Nguyen (2024).
+For validation, we adopt the empirical design of cohen2024moving.
 We construct equal-weighted quintile portfolios sorted on the moving targets score in a calendar-time framework, and we estimate Fama–MacBeth cross-sectional regressions with standard controls.
@@
 In particular, industry- and sector-level classification, combined with contextual prompt engineering, could yield even more precise and explainable measures of managerial emphasis across heterogeneous business environments.
+Figure 3: Representative example of targets that are extracted by NER-based method showing noise in the extracted targets, excerpted from Apple Inc.
+2010 Q2 earnings call transcript.
+The NER-based method extracts generic terms such as year and units as targets illustrating the noisy nature of the extraction.
+Figure 4: Representative example of targets that are extracted by NER-based method showing loss of contextual information, excerpted from NVIDIA Corporation 2025 Q2 earnings call transcript.
+The NER-based method extracts words like revenue as targets but drops surrounding qualifiers (e.g., Data Center revenue, Compute revenue), losing the specific type of revenue or operational driver.
+In addition, key product-related terms such as NVIDIA Hopper, and GPU computing are excluded from the target set even though the original approach is designed to include product entities, further underscoring the limitation of the NER-based extraction method.
+Figure 5: Representative example of targets that are extracted by LLM-based method, excerpted from NVIDIA Corporation 2025 Q2 earnings call transcript.
+Here, we highlight both the phrase in the transcript where each target is detected and the corresponding extracted target itself.
+Unlike the NER-based approach, which often extracts overly generic terms such as year and units, the LLM-based method captures richer and more semantically precise business metrics (e.g., Quarterly Data Center revenue, Quarterly Compute revenue growth), reducing the noise in the extracted set.
+Table 1: Statistics of the top 20 most frequent missing targets extracted by the NER-based method.
+The high frequency of generic phrases (e.g., the %, a % increase, the quarter) highlights the noise in the extracted targets, where context-free fragments dominate over concrete business metrics like specific revenue categories or operational drivers.
+Rank Target Freq.
+1 the % 4460
+2 a % increase 3845
+3 the range 3474
+4 % growth 3456
+5 q3 3006
+6 q4 2853
+7 the % range 2801
+8 year 2716
+9 the quarter 2499
+10 share 2400
+Rank Target Freq.
+11 dollars 2268
+12 a % 2146
+13 earnings 2115
+14 growth 2012
+15 that % 1902
+16 revenue 1882
+17 a increase 1879
+18 a range 1850
+19 q1 1555
+20 sales 1537
+Table 2: Statistics of the top 20 most frequent missing targets extracted by the LLM-based method,.
+Unlike the NER-based approach—which often produces noisy, context-free fragments—our LLM-based extraction identifies substantive business metrics (e.g., market share, revenue growth, cash flow) and quantifies when these key targets are dropped or strategically shifted across consecutive year–quarters.
+Rank Target Freq.
+1 market share 706
+2 share repurchases 589
+3 revenue growth 582
+4 pricing 562
+5 cash flow 535
+6 quarterly revenue growth 517
+7 yearly guidance 484
+8 free cash flow 480
+9 dividend 458
+10 quarterly earnings per share 455
+Rank Target Freq.
+11 quarterly share repurchases 441
+12 quarterly revenue 436
+13 profitability 435
+14 dividends 416
+15 gross margin 413
+16 operating leverage 402
+17 margins 402
+18 revenue 388
+19 share buybacks 377
+20 margin expansion 368
+Table 3: Calendar-time portfolio returns based on Moving Targets.
+We form equal-weighted quintile portfolios using the prior year’s distribution of Moving Targets across all stocks.
+Stocks enter portfolios in the month following each earnings conference call and are held until the next call.
+Panel A reports results using the original NER-based extraction method (cohen2024moving), while Panel B reports results using the LLM-based extraction method.
+We report Excess Returns (return minus risk-free rate), Fama–French 3-factor Alphas (market, size, and value), and 5-factor Alphas (market, size, value, momentum, and liquidity).
+$t$-statistics are in parentheses.
+, ∗∗, and ∗∗∗ indicate significance at the 10%, 5%, and 1% levels, respectively.
+Panel A: NER-based extraction method
+Q1 Q2 Q3 Q4 Q5 Q5–Q1
+Excess Return 0.0156∗∗∗ 0.0131∗∗∗ 0.0111∗∗∗ 0.0132∗∗∗ 0.0135∗∗∗ -0.0031
+(4.38) (3.85) (3.19) (3.99) (3.94) (-1.48)
+3-Factor Alpha 0.0047∗∗∗ 0.0027∗∗ 0.0017 0.0040∗∗∗ 0.0039∗∗ -0.0018
+(3.43) (2.12) (1.15) (2.88) (2.47) (-0.89)
+5-Factor Alpha 0.0045∗∗∗ 0.0028∗∗ 0.0016 0.0040∗∗∗ 0.0041∗∗ -0.0014
+(3.26) (2.19) (1.07) (2.79) (2.54) (-0.70)
+Panel B: LLM-based extraction method
+Q1 Q2 Q3 Q4 Q5 Q5–Q1
+Excess Return 0.0151∗∗∗ 0.0135∗∗∗ 0.0123∗∗∗ 0.0128∗∗∗ 0.0122∗∗∗ -0.0039
+(4.75) (3.94) (3.67) (3.50) (3.59) (-1.93)
+3-Factor Alpha 0.0058∗∗∗ 0.0043∗∗∗ 0.0023∗∗ 0.0022 0.0021 -0.0047∗∗
+(4.08) (3.06) (1.97) (1.44) (1.39) (-2.32)
+5-Factor Alpha 0.0058∗∗∗ 0.0045∗∗∗ 0.0027∗∗ 0.0021 0.0017 -0.0050∗∗
+(3.99) (3.10) (2.25) (1.36) (1.16) (-2.43)
+Table 4: Fama–MacBeth cross-sectional regressions of individual firm-level stock returns on Moving Targets and standard return predictors.
+We compare results using the original NER-based extraction method (cohen2024moving) and the LLM-based extraction method.
+Log(Size) is the log of market value of equity, Log(BM) is the log of book value of equity over market value of equity, Ret(-1,0) is the previous month’s return, and Ret(-12,-1) is the cumulative return from month -12 to month -1.
+$t$-statistics are shown in parentheses.
+, ∗∗, and ∗∗∗ indicate significance at the 10%, 5%, and 1% levels, respectively.
+(1) (2)
+NER-based method LLM-based method
+Ret
+Moving Targets 0.0107 -0.0465
+(1.10) (-1.70)
+Log(Size) -0.0015 -0.0029
+(-0.42) (-0.68)
+Log(BM) -0.0016 -0.0006
+(-0.36) (-0.17)
+Ret(-1, 0) 0.0404 -0.0376
+(0.98) (-0.71)
+Ret(-12,-1) -0.0019 0.0029
+(-0.22) (0.40)
+Constant 0.0129 0.0828
+(0.29) (1.43)
+R-squared 0.3897 0.3884
+N 5,615 5,615
+Table 5: LLM prompt templates used in our method to extraction of targets.
+Prompt <task> From the provided inputs, extract all performance metric and target labels and classify them into two separate lists according to the section in which they were mentioned: - presentation: items mentioned in the prepared presentation or opening remarks - analyst_qa: items mentioned during analyst question and answer sessions The term "performance metric and target labels" includes every explicitly stated financial result, non-financial result, and key business indicator discussed or mentioned.
+Both GAAP and non-GAAP labels if explicitly distinguished.
+## Section classification - presentation: items mentioned in prepared remarks/opening statements before Q&A begins.
+Typical speakers: executives with "- Executives" in the name, before the first analyst question.
+- analyst_qa: items mentioned during the Q&A (analyst questions and executive answers after Q&A starts).
+Q&A typically starts after an Operator prompt like "[Operator Instructions]" or when an analyst first speaks (speaker label contains "- Analysts") and continues until the call ends.
+- Operator-only lines do not themselves contain items, but they can mark Q&A start.
+- Include only labels explicitly stated in content - Deduplicate near-duplicates to one normalized target per section </task> <inputs>earnings-call transcript as indexed JSON dialog</inputs> <output> - target: - Short noun phrase only - **Quarter results:** prefix with "Quarterly" and do NOT include any years or quarter numbers - **Fiscal-year results:** prefix with "Yearly" and do NOT include any years - **Guidance/targets/other indicators:** no period prefixes unless inherently part of the label; avoid dates, years, and quarter numbers - No numbers, units, currency symbols, or percent signs anywhere - Prefer consistent terminology across products/segments - index: integer index of the utterance containing the quote </output> <format> {{ "presentation": List[Dict], "analyst_qa": List[Dict] }} The dictionary must follow the following structure: {{ "target": str # Normalized short label for the metric/target according to the rules "index": int # Zero-based index of the utterance in the transcript where the quote appears }} </format>